OpenStudy (cruffo):

Suppose \[S_1 = \rm span\left\{ \left|\matrix{1&0\\0&1\\}\right|, \left|\matrix{0&1\\1&0\\}\right|, \left|\matrix{1&-1\\0&1\\}\right| \right\}\] and \[S_2 = \rm span\left\{ \left|\matrix{1&0\\0&1\\}\right|, \left|\matrix{0&1\\1&0\\}\right|, \left|\matrix{0&0\\1&0\\}\right| \right\}\] Does \(S_1 = S_2\)?

OpenStudy (cruffo):

I know that \[\left|\matrix{1&0\\0&1\\}\right| - \left|\matrix{0&1\\1&0\\}\right| + \left|\matrix{0&0\\1&0\\}\right| = \left|\matrix{1&-1\\0&1\\}\right|\]

OpenStudy (cruffo):

Thing is, do I get more (or less) with \(S_2\) than I get with \(S_1\)?

OpenStudy (anonymous):